BJ can overhaul a boat's diesel inboard engine in 20 hours. His apprentice takes 60 hours to do the same job. How long would it take them working together assuming no gain or loss in efficiency?

Answer by stanbon(75887)

BJ can overhaul a boat's diesel inboard engine in 20 hours. His apprentice takes 60 hours to do the same job. How long would it take them working together assuming no gain or loss in efficiency?
---
BJ rate: 1/20 job/hr
----
App rate: 1/60 job/hr
---
Together rate: 1/x job/hr
-----
Equation:
rate + rate = together rate
1/20 + 1/60 = 1/x
Multiply thru by 60x
3x + x = 60
4x = 60
x = 15 hrs (time to do the job together)
===========
Cheers,
Stan H.